OpenStudy (anonymous):

A differentiable function f(x) is defined such that, for all values of x in its domain, (i will draw the equation). (a) what is the domain. (b) for what values of x is f(x)=3? (c)show that f'(x)=3^(2)f(x) (d)solve the differential equation in (c) to find f(x) in terms of x only.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[f(x)=3+\int_8^{x^3}f(\sqrt[3]{t})dt\] like that ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

looks to me like the domian is all real numbers

OpenStudy (anonymous):

i got that part but rest of the questions are confusing.lol

OpenStudy (anonymous):

also i am thinking it should be \[F(x)=3+\int_8^{x^3}f(\sqrt[3]{t})dt\]

OpenStudy (anonymous):

since it is clearly not the same F and f

OpenStudy (anonymous):

both say f

OpenStudy (anonymous):

\[F(x)=3\implies \int_8^{x^3}f(\sqrt[3]{x})dt=0\]

OpenStudy (anonymous):

integral is zero when upper and lower limits of integration are the same, so you have \[\int_8^{x^3}f(\sqrt[3]{t})dt=\int_8^8f(\sqrt[3]{t})dt\] and so \[x^3=8\] thefore \(x=2\)

OpenStudy (anonymous):

since the derivative of the integral is the integrand, you have \[f'(x)=f(\sqrt[3]{x^3})\times 3x^2=3x^2f(x)\]

OpenStudy (anonymous):

so i am hoping there is a typo above and you meant \[ f'(x)=3x^2f(x)\]not \[ f'(x)=3^2f(x)\]

OpenStudy (anonymous):

yes it was a typo sorry i didnt catch it.

OpenStudy (anonymous):

but it is clear yes?

OpenStudy (anonymous):

yes it is thank you.

OpenStudy (anonymous):

does that mean that d is f(x)=3ex^3-8

OpenStudy (anonymous):

ooh lets see we didn't finish

OpenStudy (anonymous):

if \(f(x)=e^{x^3}\) then \[f'(x)=3x^2e^{x^3}\] so i don't think you need the additional 3 out front

OpenStudy (anonymous):

oh damn we also have to make sure that \(f(2)=3\) don't we

OpenStudy (anonymous):

maybe i am confused, but i think we have \(f(x)=e^{x^3}+C\) and since \(f(2)=3\) we should solve \[3=e^8+C\] for C to get \(C=3-e^8\) but i could be wrong

OpenStudy (anonymous):

\[3ex^{3}-8\]thats why i thought it would be